OpenStudy (anonymous):

Find the particular solution y=f(x) to the differential equation y'=1+y/x with the initial condition f(-1)=1 I would appreciate any help

OpenStudy (anonymous):

separate variables: dy/dx(1/(1+y))=1/x dy/(1+y)=dx/x) integrate bouth sides: Ln|1+y|=Ln|x| + LnC 1+y=Cx y=Cx-1 now use initial condition: 1=C(-1)-1 from here C=-2 so your solution is: y=-2x-1

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

How would I go about finding domain from here ?